# Optimizing Large Language Models for OpenAPI Code Completion

### v1

@@ III-B Code Llama Model
This code generation model is based on another widely adopted natural language model from the same company – Llama 2. The model was trained for code infilling on an additional 500 billion tokens of code. Figure 2 outlines the training pipeline of the Code Llama model family. The model was further trained for Python specialization and fine-tuned for instruction execution, but for the sake of this research, the base Code Llama model is used.
Figure 2: Code Llama model family training pipeline. Adapter from [4].
The model is available in three sizes: 7 billion parameters, 13 billion parameters, and 34 billion parameters. In this research, the 7 billion and 13 billion parameter models are evaluated. The 34 billion parameter model was not selected for evaluation due to the high computational cost and the lack of evidence that the model would demonstrate significantly higher performance than the 13 billion parameter model. Moreover, the largest model was not trained on code infilling tasks, which is required for real-world code completion scenarios.
@@
Following the training methods originally described in the work of Bavarian et al. [57], the Code Llama model supports two prompt formats for code infilling: Prefix Suffix Middle (PSM) and Suffix Prefix Middle (SPM). The study has shown that the SPM format produces slightly better results even when the model is trained on the PSM format. The SPM format is also expected to result in a higher key-value cache hit rate as the suffix is less likely to change in a real-world code infilling scenario. Both formats are evaluated in this research.
The Code Llama model introduces four special tokens to the existing Llama 2 tokenizer: <<<PRE>>>, <<<SUF>>>, <<<MID>>>, <<<EOT>>>. Using these tokens, the PSM prompt format is defined as
@@
where the Enc()($\cdot$)( ⋅ ) is the encoding operation and $\circ$ is a token-level concatenation. The SPM format is defined as
@@
III-C Prompt Engineering Techniques
As the expected OpenAPI definition size is significantly larger than the maximum context size supported by Code Llama, it is not possible to fit the whole definition in the prompt. In this research, the impact of multiple variables on the completion performance is evaluated with the aim of finding the optimal prompt format for a real-world OpenAPI completion task.
@@ III-C Prompt Engineering Techniques
The code infilling mode is expected to be more suitable for the OpenAPI completion task as it allows the model to consider the context both before and after the cursor position. The part of the definition that goes after the cursor should allow the model to take into account components such as schemas, requests, and responses that are commonly placed at the end of the document. The suffix is also important for the model to generate syntactically correct definitions that are integral to the rest of the document.
@@ III-D Code Llama Model Fine-tuning for OpenAPI Completion
The OpenAPI completion performance of the model can be further improved by fine-tuning it on task-specific data. This section describes the dataset used for fine-tuning the 7 billion parameter version of the Code Llama model as well as the full details of the fine-tuning process.
@@
The dataset for fine-tuning was collected from the APIs-guru OpenAPI definitions directory [55]. The directory contains more than 4,000 definitions in yaml format. Analysis of the repository revealed that about 75% of the definitions in the directory are produced by a handful of major companies like Amazon, Google, and Microsoft. To avoid the dataset bias towards a specific producer, the maximum number of definitions from a single producer was limited to 20. Multiple versions of the same API were also excluded from the dataset as they are likely to contain similar definitions. Definitions from producers used in the evaluation dataset were also excluded from the fine-tuning dataset to avoid data leakage. The resulting dataset contains 990 definitions and can be accessed at the HuggingFace Hub at https://huggingface.co/datasets/BohdanPetryshyn/openapi-completion-refined.
The authors of the Code Llama model don’t reveal the details of the dataset used for training. However, considering its size (500 billion tokens), it would be naive to assume that the publicly available OpenAPI definitions were not used for training. At the same time, the Code Llama model authors claim that fine-tuning the model on already-seen data can still lead to significant performance improvements. This claim is supported by the results of the Python-specialized Code Llama version which was fine-tuned from the base Code Llama model using the subset of the original training dataset.
@@
In order to avoid the objective mismatch between the training and fine-tuning processes, the infilling prompt format was reconstructed to mirror the format described by the Code Llama model authors. Nevertheless, the specifics of the OpenAPI format still had to be taken into account. Similarly to the Code Llama training, the prefix-middle-suffix split was performed at the character level with the split positions selected randomly. The same tokens (<<<PRE>>>, <<<SUF>>>, <<<MID>>>, <<<EOT>>>) were used to format the prompt in the PSM and SPM formats and the leading spaces implicitly generated by the SentencePiece tokenizer were suppressed. However, unlike the Code Llama training, the split was performed at the context level and was not limited to documents that fit into the context size as most of the OpenAPI definitions are much larger than the context size used during fine-tuning. In the SPM format, the prefix and middle parts were concatenated after the tokenization to improve the model performance with split tokens which are expected to be encountered often in the real-world OpenAPI completion scenarios.
@@
Another consequence of the large size of the OpenAPI definitions is the bias towards full-context-length completions. As can be seen in Figure 3, the distribution of the training sample lengths is skewed toward the maximum context size of 5,120 tokens when the normal context-level split is used. This bias can cause the model to show optimal performance only when prompted with the same context size as during fine-tuning. To mitigate the bias, the input documents were split into smaller parts of random length between 4 096 and 6 144 tokens (0.8-1.2 of the 5 120 tokens context size) before the context-level split was performed. This technique is further referred to as document splitting. When used before the infilling permutation, document splitting should not be distinguishable from using full documents of uniform length distribution as the truncated prefixes and suffixes correspond to a normal real-world code completion scenario when the context size is limited and doesn’t cover the entire document.
@@ IV Results
Following the prompt engineering and fine-tuning methods described in the previous section, the Code Llama model was optimized for the OpenAPI completion task. This section presents the results of the evaluation of the optimized model and compares it to the performance of GitHub Copilot. For all the evaulations described in this section, the same set of masked OpenAPI definitions was used ensuring consistent evaluation conditions.
@@ IV-B Original Code Llama Model for OpenAPI Completion
Using the benchmark proposed in the previous chapter, an extensive set of experiments was performed to analyze the impact of a variety of variables and prompt-engineering techniques on the performance of the Code Llama model in the OpenAPI completion task. The experiments were designed to analyze the impact of a single variable at a time. When not stated otherwise, the 7 billion parameter version of the Code Llama model, the PSM prompt format (1), the 50/50 prefix-to-suffix ratio, and the 4 096 tokens context size were used.
The Code Llama models were evaluated at the Hugging Face Inference Endpoints platform [58] without quantization using a single Nvidia A10G GPU for the 7 billion parameter model and four Nvidia A10G GPUs for the 13 billion parameter model.
@@ IV-B1 Optimal prefix-to-suffix ratio
A range of prefix-to-suffix ratios from 30/70 up to 90/10 was evaluated to find the optimal proportion. The evaluation was performed on the 7 billion parameter version of the Code Llama model with a constant context size of 4 096 tokens. The optimal prefix-to-suffix ratio can depend on the mentioned factors but for the purpose of parameter space reduction, the dependencies are intentionally disregarded in this research.
@@
Figure 4: OpenAPI completion performance of the original Code Llama 7B model at a range of prefix-to-suffix ratios with a fixed context size of 4096 tokens.
@@ IV-B2 Optimal context size
Context size is another factor that affects not only the solution performance but also the speed of code generation and, consequently, throughput and resource efficiency of a real-world implementation. A range of context sizes from 1 024 to 7 168 tokens was evaluated with a constant prefix-to-suffix ratio of 50/50 and the 7 billion parameter version of the Code Llama model.
@@
Figure 5: OpenAPI completion performance of the original Code Llama 7B model at a range of context sizes with a fixed preffix-to-suffix-ratio of 50/50.
@@ IV-B4 OpenAPI metadata in prompt
The experiment was performed on the 7 billion parameter Code Llama model with a 60/40 prefix-to-suffix ratio and a range of context sizes from 4 096 to 7 168 tokens. The extended prefix size was expected to compensate for the space taken by the components section in the prompt.
@@
The obtained results might mean that the components are not usually required for correct OpenAPI completion. The model can deduct the correct component names from just the surrounding context. When used with a relatively small context size, the suggested method can even be harmful to the solution performance as it consumes space that could otherwise be taken by more valuable information in the prefix.
@@ IV-B5 Optimal model size
The Code Llama model is available in three sizes: 7 billion parameters, 13 billion parameters, and 34 billion parameters. The impact of the model size on the solution performance is evaluated by comparing the performance of the 7 billion parameter model with the performance of the 13 billion parameter model. Due to resource constraints and the inability of the 34 billion parameter model to work in the infilling mode, this model size is not evaluated in the experiment.
@@
As can be seen in Figure 8, the bigger 13 billion parameter model outperforms the smaller model at all context sizes with the only exception being correctness after 3 072 tokens. At the context size of 4 096 tokens, Code Llama 13B surpasses the results obtained with Code Llama 7B by only 1% in validity while falling behind in correctness by 5%.
Figure 8: OpenAPI completion performance of the 7 and 13 billion parameter versions of the original Code Llama model.
@@
Figure 9: Generation speed of the 7 and 13 billion parameter versions of the original Code Llama model.
@@ IV-C1 Benefits of using a mixture of PSM and SPM infilling formats
The study of Bavarian et al. suggests that using a mix of SPM and PSM formats during training leads to an increase in the infilling performance [57]. At the same time, the original Code Llama model has a known problem with infilling in the SPM format IV-B3. It is worth evaluating the efficiency of including this format in the fine-tuning process. This section compares the performance of the 7 billion parameter Code Llama model fine-tuned solely with the PSM format and with a mixture of PSM and SPM formats (joint fine-tuning). When used, the SPM format was applied to half of the training samples. Context size of 4 096 tokens was used during fine-tuning.
@@
IV-C2 Full-context-length infilling bias and document splitting technique
As revealed in the previous section, the fine-tuned model underperforms when prompted with context sizes smaller than used during fine-tuning. This can be caused by the bias towards full-context-length completions discussed in Section III-D. This section further evaluates the impact of the fine-tuning context size as well as the suggested document splitting technique on the performance of the fine-tuned model.
As can be seen in Figure 11, both 4 096 tokens and 5 120 tokens fine-tuned models underperform compared to the original Code Llama model at context sizes smaller than the corresponding fine-tuning context sizes. After this point, the performance of the fine-tuned models stabilizes. Detailed inspection of the infilling results shows that the fine-tuned models tend to generate completions that are too long and contain irrelevant information when prompted with smaller context sizes.
@@ IV-C2 Full-context-length infilling bias and document splitting technique
The model fine-tuned with the document splitting technique demonstrates more equalized performance across the context sizes. The correctness and validity of the completions generated by this model surpass the original Code Llama model at almost all context sizes with a peak increase of 20% at 6 144 and 7 168 tokens and a marginal decrease of 5.8% at 3 072 tokens. The model slightly underperforms compared to the model fine-tuned with the normal context-level split technique at the context size of 4 096 tokens when prompted with context sizes larger than 4 096 tokens, but it outperforms the same model at smaller context sizes by a significant margin.
@@ V Findings and Implications
Through the course of this research, the impact of multiple variables, prompt engineering, and fine-tuning techniques on the OpenAPI completion performance of the Code Llama model was evaluated. The most notable experimental results are summarized and compared with one of the most widely adopted commercial solutions – GitHub Copilot in Table II.
@@ V-1 Original Code Llama model is capable of outperforming GitHub Copilot
The original Code Llama 7B model, when used with the optimal prefix-to-suffix ratio of 50/50, the PSM prompt format, and the context size of 4 096 tokens, outperforms GitHub Copilot by 24.1% in correctness, generating correct OpenAPI completions 36% of the time. At the same time, the model produces valid completions 63% of the time which is a 7.3% decrease compared to Copilot. It’s worth pointing out that the GPT-3 model, which is the foundational model used by GitHub Copilot, uses 175 billion parameters which is 25 times more than in the case of the Code Llama model.
@@ V-2 Larger model does not always show significant performance improvement
The larger 13 billion parameter version of the Code Llama model demonstrated only a marginal correctness improvement compared to the smaller 7 billion parameter version. This might mean that the OpenAPI format is not complex enough to leverage the additional model capacity. Considering the quadrupled computational cost, this draws the larger version of the model not practical for use in OpenAPI completion.

### v2

@@ III-B Code Llama Model
This code generation model is based on another widely adopted natural language model from the same company – Llama 2. The model was trained for code infilling on an additional 500 billion tokens of code. Figure 2 outlines the training pipeline of the Code Llama model family. The model was further trained for Python specialization and fine-tuned for instruction execution, but for the sake of this research, the base Code Llama model is used. The 70-billion-parameter version of the model was trained and released months after the original Code Llama family and might follow a slightly different training pipeline.
Figure 2: Code Llama model family training pipeline, adapted from [4].
After the recent release of the biggest version, the model is now available in four sizes: 7, 13, 34, and 70 billion parameters. This study is focused on the smaller 7 and 13-billion-parameter versions of the model because the bigger 34-billion-parameter version does not support the infilling mode of operation ($\leftrightarrow$” in Figure 2, discussed below) and the computational requirements of the 70-billion-parameter version exceed the capabilities of the evaluation platform.
@@
Following the training methods originally described in the work of Bavarian et al. [27], the Code Llama model supports two prompt formats for code infilling: Prefix Suffix Middle (PSM) and Suffix Prefix Middle (SPM). The study has shown that the SPM format produces slightly better results even when the model is trained on the PSM format. The SPM format is also expected to result in a higher key-value cache hit rate as the suffix is less likely to change in a real-world code completion scenario. Both formats are evaluated in this research.
The Code Llama model introduces four special tokens to the existing Llama 2 tokenizer: <<<PRE>>>, <<<SUF>>>, <<<MID>>>, <<<EOT>>>. Using these tokens, the PSM prompt format is defined in [27] as
@@
where the Enc()($\cdot$)( ⋅ ) is the encoding operation and $\circ$ is a token-level concatenation. The SPM format is defined in [27] as
@@
III-C Prompt Engineering Techniques for OpenAPI Completion
As the expected OpenAPI definition size is significantly larger than the maximum context size supported by Code Llama, it is not possible to fit the whole definition in the prompt. In this research, the impact of multiple variables on the completion performance is evaluated to find the optimal prompt format for a real-world OpenAPI completion task.
Code Llama’s infilling mode is more suitable for the OpenAPI completion task than the autoregressive generation mode as it allows the model to consider the context both before and after the cursor position. The part of the definition that goes after the cursor should allow the model to take into account components such as schemas, requests, and responses that are commonly placed at the end of the document. The suffix is also important for the model to generate syntactically correct definitions that are integral to the rest of the document. Editing at the end of a document can be considered an infilling task where the suffix is empty. In the real world, API designers are more likely to edit existing definitions by adding new operations or parameters in the middle of the document rather than at the end of it.
@@ III-D Code Llama Model Fine-tuning for OpenAPI Completion
The OpenAPI completion performance of the model can be further improved by fine-tuning it on task-specific data. This section describes the dataset used for fine-tuning the 7-billion-parameter version of the Code Llama model as well as the full details of the fine-tuning process.
@@
The dataset for fine-tuning was collected from the APIs-guru OpenAPI definitions directory [56]. The directory contains more than 4 000 definitions in YAML format. Analysis of the repository revealed that about 75% of the definitions in the directory are produced by a handful of major companies like Amazon, Google, and Microsoft. To avoid the dataset bias towards a specific producer, the maximum number of definitions from a single producer was limited to 20. Multiple versions of the same API were also excluded from the dataset as they are likely to contain similar definitions. Definitions from producers used in the evaluation dataset were also excluded from the fine-tuning dataset to avoid data leakage. The resulting dataset contains 990 definitions and is accessible at the HuggingFace Hub at https://huggingface.co/datasets/BohdanPetryshyn/openapi-completion-refined.
@@
To avoid the objective mismatch between the training and fine-tuning processes, the infilling prompt format was reconstructed to mirror the format described by the Code Llama model authors. Nevertheless, the specifics of the OpenAPI format had to be taken into account. Similarly to the Code Llama training, the prefix-middle-suffix split was performed at the character level with the split positions selected randomly. The same tokens (<<<PRE>>>, <<<SUF>>>, <<<MID>>>, <<<EOT>>>) were used to format the prompt in the PSM (1) and SPM (2) formats and the leading spaces implicitly generated by the SentencePiece tokenizer were suppressed. However, unlike the Code Llama training, the split was performed at the context level and was not limited to documents that fit into the context size as most of the OpenAPI definitions are much larger than the context size used during fine-tuning. In the SPM format, the prefix and middle parts were concatenated after the tokenization to improve the model performance with split tokens which are expected to be encountered often in the real-world OpenAPI completion scenarios.
@@
Another consequence of the large size of the OpenAPI definitions is the bias towards training-time context size infilling that might be developed when fine-tuning the model using the conventional FIM training objective as described in the work of Bavarian et. al. [27]. This bias can cause the model to show optimal performance only when prompted with the same context size as during fine-tuning. As can be seen in Figure 3, the distribution of the training sample lengths is skewed toward the maximum context size of 5 120 tokens when the conventional context-level FIM transformation is used. To mitigate the bias, the input documents were split into smaller pieces of random length between 4 096 and 6 144 tokens (0.8-1.2 of the 5 120 tokens context size) before proceeding to the context-level FIM transformation. This made sure the context chunks included multiple documents of various lengths more often and evened the distribution of the training sample lengths. This technique is further referred to as document splitting. When used before the FIM transformation, document splitting should not be distinguishable from using full documents of uniform length distribution as the truncated prefixes and suffixes correspond to a normal real-world code completion scenario when the context size is limited and does not cover the entire document.
@@ IV Results
Following the prompt engineering and fine-tuning methods described in the previous section, the Code Llama model was optimized for the OpenAPI completion task. This section presents the results of the evaluation of the optimized model and compares it to the performance of GitHub Copilot. For all the evaluations described in this section, the same set of masked OpenAPI definitions was used ensuring consistent evaluation conditions.
@@ IV-B Original Code Llama Model for OpenAPI Completion
Using the benchmark proposed in the previous chapter, an extensive set of experiments was performed to analyze the impact of a variety of variables and prompt-engineering techniques on the performance of the Code Llama model in the OpenAPI completion task. The experiments were designed to analyze the impact of a single variable at a time. When not stated otherwise, the 7-billion-parameter version of the Code Llama model, the PSM prompt format (1), the 50/50 prefix-to-suffix ratio, and the 4 096 tokens context size were used.
The Code Llama models were evaluated at the Hugging Face Inference Endpoints platform [58] using the simple greedy search decoding strategy. Without quantization, a single Nvidia A10G GPU was used for the 7-billion-parameter model. For the 13-billion-parameter model, four Nvidia A10G GPUs were used.
@@ IV-B1 Optimal prefix-to-suffix ratio
A range of prefix-to-suffix ratios from 30/70 up to 90/10 was evaluated to find the optimal proportion. The evaluation was performed on the 7-billion-parameter version of the Code Llama model with a constant context size of 4 096 tokens. The optimal prefix-to-suffix ratio can depend on the above-mentioned factors but for parameter space reduction, the dependencies are intentionally disregarded in this research.
@@
Figure 4: OpenAPI completion performance of the original Code Llama 7B model for different prefix-to-suffix ratios with a fixed context size of 4 096 tokens.
@@ IV-B2 Optimal context size
Context size is another factor that affects not only the solution performance but also the speed of code generation and, consequently, throughput and resource efficiency of a real-world implementation. A range of context sizes from 1 024 to 7 168 tokens was evaluated with a constant prefix-to-suffix ratio of 50/50 and the 7-billion-parameter version of the Code Llama model.
@@
Figure 5: OpenAPI completion performance of the original Code Llama 7B model for different context sizes with a fixed preffix-to-suffix-ratio of 50/50.
@@ IV-B4 OpenAPI metadata in prompt
This experiment evaluates the impact of including the components section of the OpenAPI definition in the prompt prefix as described in Section III-C. The experiment was performed on the 7-billion-parameter Code Llama model with a 60/40 prefix-to-suffix ratio and a range of context sizes from 4 096 to 7 168 tokens. The extended prefix size was expected to compensate for the space taken by the components section in the prompt.
@@
The obtained results might mean that the components are not necessarily required for correct OpenAPI completion. The model can deduct the correct component names from just the surrounding context. When used with a relatively small context size, the suggested method can even be harmful to the solution performance as it consumes space that could otherwise be taken by more valuable information in the prefix.
@@ IV-B5 Optimal model size
The Code Llama model is available in four sizes: 7, 13, 34, and 70 billion parameters. The impact of the model size on the solution performance is evaluated by comparing the performance of the 7 and 13-billion-parameter models. Due to the resource constraints and the limitations of the larger models discussed in Section III-B, the 34 and 70-billion-parameter models were not evaluated.
@@
As can be seen in Figure 8, the bigger 13-billion-parameter model outperforms the smaller model at all context sizes with the only exception being correctness after 3 072 tokens. At the context size of 4 096 tokens, Code Llama 13B surpasses the results obtained with Code Llama 7B by only 1% in validity while falling behind in correctness by 5%.
Figure 8: OpenAPI completion performance of the 7 and 13-billion-parameter versions of the original Code Llama model.
@@
Figure 9: Generation speed of the 7 and 13-billion-parameter versions of the original Code Llama model.
@@ IV-C1 Benefits of using a mixture of PSM and SPM infilling formats
The study of Bavarian et al. suggests that using a mix of SPM and PSM formats during training leads to an increase in the infilling performance [27]. At the same time, the original Code Llama model has a known problem with infilling in the SPM format IV-B3. It is worth evaluating the efficiency of including this format in the fine-tuning process. This section compares the performance of the 7-billion-parameter Code Llama model fine-tuned solely with the PSM format and with a mixture of PSM and SPM formats (joint fine-tuning). When used, the SPM format was applied to half of the training samples. Context size of 4 096 tokens and a 50/50 prefix-suffix split was used during fine-tuning.
@@
IV-C2 Training-time context size infilling bias and document splitting technique
As revealed in the previous section, the fine-tuned model underperforms when prompted with context sizes smaller than used during fine-tuning (Figure 10). This can be caused by the bias towards training-time context size infilling discussed in Section III-D. This section further evaluates the impact of the fine-tuning context size as well as the suggested document splitting technique on the performance of the fine-tuned model.
As can be seen in Figure 11, the fine-tuned models with both 4 096 tokens and 5 120 tokens underperform compared to the original Code Llama model at context sizes smaller than the corresponding fine-tuning context sizes. After this point, the performance of the fine-tuned models stabilizes. Detailed inspection of the infilling results shows that the fine-tuned models tend to generate completions that are too long and contain irrelevant information when prompted with smaller context sizes.
@@ IV-C2 Training-time context size infilling bias and document splitting technique
The model fine-tuned with the document splitting technique demonstrates more equalized performance across the context sizes. The correctness and validity of the completions generated by this model surpass the original Code Llama model at almost all context sizes with a peak increase of 20% at 6 144 and 7 168 tokens and a marginal decrease of 5.8% at 3 072 tokens. The model slightly underperforms compared to the model fine-tuned with the conventional context-level FIM technique at the context size of 4 096 tokens when prompted with context sizes larger than 4 096 tokens, but it outperforms the same model at smaller context sizes by a significant margin.
@@ V Findings and Implications
Through the course of this research, the impact of multiple variables, prompt engineering, and fine-tuning techniques on the OpenAPI completion performance of the Code Llama model was evaluated. The most notable experiment results are summarized and compared in Table II.
@@ V-1 Original Code Llama model is capable of outperforming GitHub Copilot
The original Code Llama 7B model, when used with the optimal prefix-to-suffix ratio of 50/50, the PSM prompt format, and the context size of 4 096 tokens, outperforms GitHub Copilot by 24.1% in correctness, generating correct OpenAPI completions 36% of the time. At the same time, the model produces valid completions 63% of the time which is a 7.3% decrease compared to Copilot. It is worth pointing out that the GPT-3 model, which is the foundational model used by GitHub Copilot, uses 175 billion parameters which is 25 times more than in the case of the Code Llama 7B model.
@@ V-2 Larger model does not always show significant performance improvement
The larger 13-billion-parameter version of the Code Llama model demonstrated only a marginal correctness improvement compared to the smaller 7-billion-parameter version. This might mean that the OpenAPI format is not complex enough to leverage the additional model capacity. Considering the quadrupled computational cost, this renders the larger version of the model not practical for use in OpenAPI completion.